A chemist needs to have a 70% acid solution. He has 5 gallons of a 90% acid solution. How much pure water must be added to the 90% acid solution to obtain the 70% acid solution he requires?
Answer by stanbon(75887)

A chemist needs to have a 70% acid solution. He has 5 gallons of a 90% acid solution. How much pure water must be added to the 90% acid solution to obtain the 70% acid solution he requires?
---------------------------
Equation:
acid + acid = acid
0.90*5 + 0*x = 0.70(5+x)
Multiply thru by 100 to get:
90*5 + 0 = 5 *70 + 70x
450 = 350 + 70x
70x = 100
x = 10/7 gallons
x = 1.43 gallons
========================
Cheers,
Stan H.